OpenStudy (anonymous):

Max has just won some money on a game show! He has the option to take a lump sum payment of $500,000 now or get paid an annuity of $4,900 per month for the next 10 years. Assuming the growth rate of the economy is 2.9% compounding annually over the next 10 years, which is the better deal for Max and by how much?

ganeshie8 (ganeshie8):

use the Present Value formula for annuity

ganeshie8 (ganeshie8):

Present Value = \(\large C\left(\dfrac{1}{i} - \dfrac{1}{i(1+i)^{nt}}\right) \)

ganeshie8 (ganeshie8):

^^this one..

ganeshie8 (ganeshie8):

its an annuity cuz he is getting money each month...

OpenStudy (anonymous):

PV= 500,000(1/.29-1/.29(1+.29)^120)

OpenStudy (anonymous):

Is that right?

ganeshie8 (ganeshie8):

Nope

OpenStudy (anonymous):

Is it all wrong? D:

ganeshie8 (ganeshie8):

you should get : Present Value = \(\large 4900*12*\left(\dfrac{1}{0.029} - \dfrac{1}{0.029(1+0.029)^{10}}\right) \)

ganeshie8 (ganeshie8):

yearly u are getting a payment of 4900*12, so, \(\large C = 4900*12\)

ganeshie8 (ganeshie8):

since this in annual compounding, \(n = 1\) So, \(\large i = \frac{r}{n} = 0.029\)

ganeshie8 (ganeshie8):

evaluate and see how much u get for Present Value of this annuity

ganeshie8 (ganeshie8):

compare it with the lump sum payment of $500,000 and decide

OpenStudy (anonymous):

670980?

ganeshie8 (ganeshie8):

nope

ganeshie8 (ganeshie8):

wolfram says, $504,145

ganeshie8 (ganeshie8):

^^

OpenStudy (anonymous):

ohh okay so a. Lump Sum: by $77,462.75 b. Lump Sum: by $4,145.41 c. Annuity: by $88,000.00 d. Annuity: by $4,145.41 it would be b then?

OpenStudy (anonymous):

no sorry it would be D

ganeshie8 (ganeshie8):

Yes ! Annuity pays more here so it would be D